dddarryl:

what mass of Fe can be obtained from 3.80 grams Fe2O3 reacting with excess CO?

justjm:

\(3.8\ g\ Fe_{2}O_{3}\cdot\frac{1\ mol\ Fe_{2}O_{3}}{159.7\ g\ Fe_{2}O_{3}}\cdot\frac{2\ mol\ Fe}{1\ mol\ Fe_{2}O_{3}}\cdot\frac{55.845}{1\ mol\ Fe}\)